morty was born 6 years berfore petunia the sum of thier age is 28.. how old are they?
Answer by JulietG(1812)

M=P+6
M+P=28
Substitute the known value of M from the first equation into the second
M+P=28
(P+6)+P=28
Add
2P+6=28
Subtract 6 from each side
2P=22
Divide each side by 2
P = 11
If Petunia is 11 and Morty was born six years earlier, then Morty is 17
17 + 11 = 28